# Imaging two-body correlations in atomic nuclei via low- and high-energy processes

### v1

@@ B.4 Second quantization formulation
$|i\rangle\equiv c^{\dagger}_{i}|0\rangle,$ (23)
where $|0\rangle$ denotes the physical vacuum and where the operator $c^{\dagger}_{i}$ creates a nucleon in state $|i\rangle$, the second-quantized form of the eccentricity operator reads as
$\mathcal{E}^{(1)}_{\ell}$ $=\sum_{ij}(\mathcal{\epsilon}^{(1)}_{\ell})_{ij}\,c^{\dagger}_{i}c_{j}\;,$ (24)
@@
$\mathcal{E}^{(2)\,1\text{b}}_{\ell}$ $\equiv\sum_{ij}(\bar{\mathcal{\epsilon}}^{(2)}_{\ell})_{ij}\,c^{\dagger}_{i}c_{j}\;,$ (26a)
$\mathcal{E}^{(2)\,2\text{b}}_{\ell}$ $\equiv\frac{1}{4}\sum_{ijkl}(\bar{\mathcal{\epsilon}}^{(2)}_{\ell})_{ijkl}\,c^{\dagger}_{i}c^{\dagger}_{j}c_{l}c_{k}\;,$ (26b)
@@
B.5 Scalar component of $\mathcal{E}^{(2)}_{\ell}$
@@
$\mathcal{E}^{(2)}_{\ell}=\sum_{L=0}^{2\ell}[\mathcal{E}^{(2)}_{\ell}]_{L0}\,$, (28)
@@
L0 $\equiv\sum_{i_{1}i_{2}=1}^{A}C^{L0}_{\ell\ell\ell-\ell}\frac{r^{\ell}_{i_{1}}r^{\ell}_{i_{2}}}{c_{\ell}c_{-\ell}}$ (29)
$\quad\quad\times\sum_{\mu=-\ell}^{\ell}C^{L0}_{\ell\mu\ell-\mu}Y^{\mu}_{\ell}(\Omega_{i_{1}})Y^{-\mu}_{\ell}(\Omega_{i_{2}})\;,$
where $C^{L0}_{l_{1}m_{1}l_{2}m_{2}}$ denotes Clebsch-Gordan (CG) coefficients. For $L=0$, CG coefficients read as
@@
such that the scalar component of the operators of multipolarity $\ell=2$ and $\ell=3$ of present interest read as
$[\mathcal{E}^{(2)}_{2}]_{00}$ $=\frac{32\pi}{75}\sum_{i_{1}i_{2}=1}^{A}r^{2}_{i_{1}}r^{2}_{i_{2}}\sum_{\mu=-2}^{2}(-1)^{\mu}Y^{\mu}_{2}(\Omega_{i_{1}})Y^{-\mu}_{2}(\Omega_{i_{2}})\;,$ (31a)
$[\mathcal{E}^{(2)}_{3}]_{00}$ $=\frac{64\pi}{245}\sum_{i_{1}i_{2}=1}^{A}r^{3}_{i_{1}}r^{3}_{i_{2}}\sum_{\mu=-3}^{3}(-1)^{\mu}Y^{\mu}_{3}(\Omega_{i_{1}})Y^{-\mu}_{3}(\Omega_{i_{2}})\;.$ (31b)
The one-body and two-body components of the operators of these two operators can be obtained following the same steps as before.
B.6 PGCM mean-square eccentricity
@@
$\rho_{s_{1}r_{1}}(p,q)$ $\equiv\frac{\langle\Phi(p)|c^{\dagger}_{r_{1}}c_{s_{1}}|\Phi(q)\rangle}{\langle\Phi(p)|\Phi(q)\rangle}\;,$ (33)
$\kappa_{s_{1}s_{2}}(p,q)$ $\equiv\frac{\langle\Phi(p)|c_{s_{2}}c_{s_{1}}|\Phi(q)\rangle}{\langle\Phi(p)|\Phi(q)\rangle}\;,$ (34)
$\bar{\kappa}^{*}_{r_{1}r_{2}}(p,q)$ $\equiv\frac{\langle\Phi(p)|c^{\dagger}_{r_{2}}c^{\dagger}_{r_{1}}|\Phi(q)\rangle}{\langle\Phi(p)|\Phi(q)\rangle}\;.$ (35)
@@
$\langle\Phi(p)|\mathcal{E}^{(2)\,2\text{b}}_{\ell}|\Phi(q)\rangle$ $=\frac{1}{4}\sum_{ijkl}(\bar{\mathcal{\epsilon}}^{(2)}_{\ell})_{ijkl}[\rho_{ki}(p,q)\rho_{lj}(p,q)-\rho_{kj}(p,q)\rho_{li}(p,q)+\bar{\kappa}^{*}_{ij}(p,q)\kappa_{kl}(p,q)]$
$=\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{\ell}\rho(p,q)]\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{-\ell}\rho(p,q)]-\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{\ell}\rho(p,q)\mathcal{\epsilon}^{(1)}_{-\ell}\rho(p,q)]$
$\quad+\frac{1}{2}\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{\ell}\kappa(p,q)(\mathcal{\epsilon}^{(1)}_{-\ell})^{T}\bar{\kappa}^{\dagger}(p,q)]-\frac{1}{2}\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{\ell}\kappa^{T}(p,q)(\mathcal{\epsilon}^{(1)}_{-\ell})^{T}\bar{\kappa}^{\dagger}(p,q)]\;,$ (37)
where advantage was taken in the last line of the separable character of the two-body matrix elements (Eq. (27b)) such that the the kernel is much less costly to compute than for a general two-body operator.
B.7 Classical rigid rotor interpretation
The interpretation of the normalized mean-squared eccentricity $\langle\epsilon^{2}_{\ell}\rangle_{0^{+}}$ within the classical rigid-rotor (RR) model allows for insightful conclusions regarding the nature of multi-particle correlations [11].
@@ B.7 Classical rigid rotor interpretation
$\rho^{(1)}_{\mathrm{intr}}(\mathbf{r}^{\prime})$ $\equiv A\frac{e^{-r^{\prime 2}/2R(\theta^{\prime})^{2}}}{(2\pi)^{3/2}R^{3}}\;,$ (38)
@@
$R(\theta^{\prime})\equiv R[1+\beta_{20}Y_{2}^{0}(\theta^{\prime})+\beta_{30}Y_{3}^{0}(\theta^{\prime})],$ (39)
and where $\theta^{\prime}$ is the angle in the intrinsic frame with respect to the symmetry axis.
@@
$\rho^{(1)}_{\mathrm{intr}}(\mathbf{r}^{\prime})$ $\approx A\frac{e^{-r^{\prime 2}/2R^{2}}}{(2\pi)^{3/2}R^{3}}[1+\frac{r^{\prime 2}}{R^{2}}\beta_{20}Y_{2}^{0}(\theta^{\prime})+\frac{r^{\prime 2}}{R^{2}}\beta_{30}Y_{3}^{0}(\theta^{\prime})]\;,$ (40)
@@
$\rho^{(1)}_{\mathrm{lab}}(\mathbf{r})\equiv\frac{1}{4\pi}\int_{\Omega_{\mathrm{or}}}\rho^{(1)}_{\mathrm{intr}}(R_{\Omega_{\rm or}}^{-1}\mathbf{r})=A\frac{e^{-r^{\prime 2}/2R^{2}}}{(2\pi)^{3/2}R^{3}}\;,$ (41)
@@
Calculating the expectation value of the two-body part of the squared eccentricity operator requires the local two-body density, which is obtained in the classical RR model according to [^5] Because of its classical character, the RR model only consider the direct part of the two-body density, which is thus not antisymmetric under the exchange of the two particles coordinates. As a result, its integration over space is equal to $A^{2}$ instead of A(A1)A(A-1) as it would be the case for a quantal two-body density.
$\rho^{(2)}_{\mathrm{lab}}(\mathbf{r}_{1},\mathbf{r}_{2})$ $\equiv\frac{1}{4\pi}\int_{\Omega_{\mathrm{or}}}\rho^{(1)}_{\mathrm{intr}}(R_{\Omega_{\rm or}}^{-1}\mathbf{r}_{1})\rho^{(1)}_{\mathrm{intr}}(R_{\Omega_{\rm or}}^{-1}\mathbf{r}_{2})$ (42)
$=A^{2}\frac{e^{-(r^{2}_{1}+r^{2}_{2})/2R^{2}}}{32\pi^{4}R^{6}}\times[4\pi$
@@
Inserting Eqs. (41) and (42) into Eq. (19) leads to
$\langle\epsilon^{2}_{2}\rangle^{\text{RR}}_{0^{+}_{1}}=\frac{1}{A}+\frac{3}{4\pi}\beta^{2}_{20}\;,\quad\langle\epsilon^{2}_{3}\rangle^{\text{RR}}_{0^{+}_{1}}=\frac{16}{3\pi A}+\frac{2048}{245\pi^{3}}\beta^{2}_{30}\;.$ (44)
@@
$\mathcal{B}^{2}_{\ell}(\text{HE})\equiv\frac{4\pi}{3}\langle\epsilon^{2}_{\ell}\rangle^{2\text{b}}\;.$ (45)
Appendix C Kumar operator
C.1 Definition

### v2

@@ II.4 Second quantization I
$|i\rangle\equiv a^{\dagger}_{i}|0\rangle,$ (25)
where $|0\rangle$ denotes the physical vacuum and where the operator $a^{\dagger}_{i}$ creates a nucleon in state $|i\rangle$, the second-quantized form of the eccentricity operator reads as
$\mathcal{E}^{(1)}_{\ell}$ $=\sum_{ij}(\mathcal{\epsilon}^{(1)}_{\ell})_{ij}\,a^{\dagger}_{i}a_{j}\;,$ (26)
@@
$\mathcal{E}^{(2)\,1\text{b}}_{\ell}$ $\equiv\sum_{ij}(\bar{\mathcal{\epsilon}}^{(2)}_{\ell})_{ij}\,a^{\dagger}_{i}a_{j}\;,$ (28a)
$\mathcal{E}^{(2)\,2\text{b}}_{\ell}$ $\equiv\frac{1}{4}\sum_{ijkl}(\bar{\mathcal{\epsilon}}^{(2)}_{\ell})_{ijkl}\,a^{\dagger}_{i}a^{\dagger}_{j}a_{l}a_{k}\;,$ (28b)
@@
II.5 Scalar component of $\mathcal{E}^{(2)}_{\ell}$
@@
$\mathcal{E}^{(2)}_{\ell}=\sum_{L=0}^{2\ell}[\mathcal{E}^{(2)}_{\ell}]_{L0}\equiv\sum_{L=0}^{2\ell}C^{L0}_{\ell\ell\ell-\ell}[\tilde{\mathcal{E}}^{(2)}_{\ell}]_{L0}\,$, (30)
@@
L0 $\equiv\sum_{i_{1}i_{2}=1}^{A}\frac{r^{\ell}_{i_{1}}r^{\ell}_{i_{2}}}{c_{\ell}c_{-\ell}}[Y^{\ell}_{\ell}(\Omega_{i_{1}})Y^{-\ell}_{\ell}(\Omega_{i_{2}})]_{L0}$ (31)
$\equiv\sum_{i_{1}i_{2}=1}^{A}\frac{r^{\ell}_{i_{1}}r^{\ell}_{i_{2}}}{c_{\ell}c_{-\ell}}\sum_{\mu=-\ell}^{\ell}C^{L0}_{\ell\mu\ell-\mu}Y^{\mu}_{\ell}(\Omega_{i_{1}})Y^{-\mu}_{\ell}(\Omega_{i_{2}})$
$\equiv\sum_{\mu=-\ell}^{\ell}C^{L0}_{\ell\mu\ell-\mu}E^{+\,(1)}_{\ell\mu}E^{-\,(1)}_{\ell-\mu}$
$\equiv[E^{+\,(1)}_{\ell}\times E^{-\,(1)}_{\ell}]_{L0}\;,$
where $C^{LM}_{l_{1}m_{1}l_{2}m_{2}}$ denotes Clebsch-Gordan (CG) coefficients. Taking $l_{1}=l_{2}=m_{1}=-m_{2}$, the CG coefficient in Eq. (30) reads as
$C^{L0}_{\ell\ell\ell-\ell}=(2\ell)!\sqrt{\frac{2L+1}{(2\ell+L+1)!(2\ell-L)!}}\;,$ (32)
The operator $\tilde{\mathcal{E}}^{(2)}_{\ell}]_{L0}$ is the sum of a one-body operator and of a two-body operator, i.e.,
$[\tilde{\mathcal{E}}^{(2)}_{\ell}]_{L0}\equiv[\tilde{\mathcal{E}}^{(2)}_{\ell}]^{1\text{b}}_{L0}+[\tilde{\mathcal{E}}^{(2)}_{\ell}]^{2\text{b}}_{L0}\;,$ (33)
with
$[\tilde{\mathcal{E}}^{(2)}_{\ell}]_{L0}^{1\mathrm{b}}$ $\equiv\sum_{i_{1}=1}^{A}\frac{r^{2\ell}_{i_{1}}}{c_{\ell}c_{-\ell}}\sum_{\mu=-\ell}^{\ell}C^{L0}_{\ell\mu\ell-\mu}Y^{\mu}_{\ell}(\Omega_{i_{1}})Y^{-\mu}_{\ell}(\Omega_{i_{1}})$
$=\sum_{i_{1}=1}^{A}\frac{r^{2\ell}_{i_{1}}}{c_{\ell}c_{-\ell}}\frac{2\ell+1}{\sqrt{4\pi(2L+1)}}C^{L0}_{\ell 0\ell 0}Y^{0}_{L}(\Omega_{i_{1}})\;,$ (34a)
$[\tilde{\mathcal{E}}^{(2)}_{\ell}]^{2\text{b}}_{L0}$ $\equiv\frac{1}{2}\sum_{i_{1}\neq i_{2}=1}^{A}2\frac{r^{\ell}_{i_{1}}r^{\ell}_{i_{2}}}{c_{\ell}c_{-\ell}}$
$\quad\quad\times\sum_{\mu=-\ell}^{\ell}C^{L0}_{\ell\mu\ell-\mu}Y^{\mu}_{\ell}(\Omega_{i_{1}})Y^{-\mu}_{\ell}(\Omega_{i_{2}})\;,$ (34b)
where the inverse Clebsch-Gordan series relation [44] was employed for the one-body term. For $L=0$, $Y_{0}^{0}(\Omega)=1/4\pi$ and the needed CG coefficients read as
@@
such that the scalar one-body and two-body components read as
$[\tilde{\mathcal{E}}^{(2)}_{\ell}]_{00}^{\mathrm{1b}}$ $=\sum_{i_{1}=1}^{A}\frac{(-1)^{\ell}}{4\pi}\frac{\sqrt{2\ell+1}}{c_{\ell}c_{-\ell}}r^{2\ell}_{i_{1}}\,$, (36a)
$[\tilde{\mathcal{E}}^{(2)}_{\ell}]_{00}^{\mathrm{2b}}$ $=\frac{1}{2}\sum_{i_{1}\neq i_{2}=1}^{A}\frac{2}{c_{\ell}c_{-\ell}}r^{\ell}_{i_{1}}r^{\ell}_{i_{2}}$
$\quad\quad\times\sum_{\mu=-\ell}^{\ell}\frac{(-1)^{\ell-\mu}}{\sqrt{2\ell+1}}Y^{\mu}_{\ell}(\Omega_{i_{1}})Y^{-\mu}_{\ell}(\Omega_{i_{2}})\,.$ (36b)
Combining Eqs. (32) and (36) allows one to express the scalar ($L=0$) one-body and two-body components of the operator $\mathcal{E}^{(2)}_{\ell}$ as
$[\mathcal{E}^{(2)}_{\ell}]_{00}^{\mathrm{1b}}$ $=\sum_{i_{1}=1}^{A}\frac{1}{4\pi}\frac{(-1)^{\ell}}{c_{\ell}c_{-\ell}}r^{2\ell}_{i_{1}}\,$, (37a)
$[\mathcal{E}^{(2)}_{\ell}]_{00}^{\mathrm{2b}}$ $=\frac{1}{2}\sum_{i_{1}\neq i_{2}=1}^{A}\frac{2}{2\ell+1}\frac{(-1)^{\ell}}{c_{\ell}c_{-\ell}}r^{\ell}_{i_{1}}r^{\ell}_{i_{2}}$
$\quad\quad\times\sum_{\mu=-\ell}^{\ell}(-1)^{\mu}Y^{\mu}_{\ell}(\Omega_{i_{1}})Y^{-\mu}_{\ell}(\Omega_{i_{2}})\,$, (37b)
where the expression of $\mathcal{E}^{(2)}_{\ell}]_{00}^{\mathrm{1b}}$ is the correct form of the erroneous Eq. (38) in the SM of Ref. [15]. Specifying the above for multipolarities $\ell=2$ and $\ell=3$ of present interest provides
$[\mathcal{E}^{(2)}_{2}]^{1\text{b}}_{00}$ $=\frac{8}{15}\sum_{i_{1}=1}^{A}r^{4}_{i_{1}}\;,$ (38a)
$[\mathcal{E}^{(2)}_{2}]^{2\text{b}}_{00}$ $=\frac{32\pi}{75}\sum_{i_{1}\neq i_{2}=1}^{A}r^{2}_{i_{1}}r^{2}_{i_{2}}\sum_{\mu=-2}^{2}(-1)^{\mu}Y^{\mu}_{2}(\Omega_{i_{1}})Y^{-\mu}_{2}(\Omega_{i_{2}})\;,$ (38b)
$[\mathcal{E}^{(2)}_{3}]^{1\text{b}}_{00}$ $=\frac{16}{35}\sum_{i_{1}=1}^{A}r^{6}_{i_{1}}\;,$ (38c)
$[\mathcal{E}^{(2)}_{3}]^{2\text{b}}_{00}$ $=\frac{64\pi}{245}\sum_{i_{1}\neq i_{2}=1}^{A}r^{3}_{i_{1}}r^{3}_{i_{2}}\sum_{\mu=-3}^{3}(-1)^{\mu}Y^{\mu}_{3}(\Omega_{i_{1}})Y^{-\mu}_{3}(\Omega_{i_{2}})\;.$ (38d)
II.6 Second quantization II
Expressing the one-body operators $E^{\pm\,(1)}_{\ell\mu}$ in their second-quantized form
$E^{\pm\,(1)}_{\ell\mu}=\sum_{ij}(e^{\pm\,(1)}_{\ell\mu})_{ij}\,a^{\dagger}_{i}a_{j}\;,$ (39)
with
$(e^{\pm\,(1)}_{\ell\mu})_{ij}\equiv\langle i|\frac{r^{\ell}}{c_{\pm\ell}}\,Y_{\ell}^{\mu}(\Omega)|j\rangle\,$, (40)
one observes that the one-body matrix elements of the eccentricity operator $(\epsilon^{(1)}_{\pm\ell})_{ij}\equiv(e^{\pm\,(1)}_{\ell\pm\ell})_{ij}$. Given Eqs. (39) and (40) the one-body and two-body components of $\tilde{\mathcal{E}}^{(2)}_{\ell}]_{L0}$ are obtained via the application of standard Wick’s theorem [46] as
$[\tilde{\mathcal{E}}^{(2)}_{\ell}]_{L0}^{1\mathrm{b}}$ $\equiv\sum_{ij}([\bar{\tilde{\mathcal{\epsilon}}}^{(2)}_{\ell}]_{L0})_{ij}\,a^{\dagger}_{i}a_{j}\;,$ (41a)
$[\tilde{\mathcal{E}}^{(2)}_{\ell}]_{L0}^{2\mathrm{b}}$ $\equiv\frac{1}{4}\sum_{ijkl}([\bar{\tilde{\mathcal{\epsilon}}}^{(2)}_{\ell}]_{L0})_{ijkl}\,a^{\dagger}_{i}a^{\dagger}_{j}a_{l}a_{k}\;,$ (41b)
with
$([\bar{\tilde{\mathcal{\epsilon}}}^{(2)}_{\ell}]_{L0})_{ij}$ $=\sum_{\mu=-\ell}^{\ell}C^{L0}_{\ell\mu\ell-\mu}\sum_{k}(e^{+\,(1)}_{\ell\mu})_{ik}(e^{-\,(1)}_{\ell-\mu})_{kj}\,$, (42a)
$([\bar{\tilde{\mathcal{\epsilon}}}^{(2)}_{\ell}]_{L0})_{ijkl}$ $=\sum_{\mu=-\ell}^{\ell}C^{L0}_{\ell\mu\ell-\mu}$
$\times\sum_{k}2[(e^{+\,(1)}_{\ell\mu})_{ik}(e^{-\,(1)}_{\ell-\mu})_{jl}-(e^{+\,(1)}_{\ell\mu})_{il}(e^{-\,(1)}_{\ell-\mu})_{jk}]\;,$ (42b)
where the two-body matrix elements are antisymmetrized. Of course, the matrix elements of the one and two-body parts of $\mathcal{E}^{(2)}_{\ell}]_{L0}$ are trivially obtained by multiplying the above with $C^{L0}_{\ell\ell\ell-\ell}$.
II.7 PGCM mean-square eccentricity
@@
$\rho_{s_{1}r_{1}}(p,q)$ $\equiv\frac{\langle\Phi(p)|a^{\dagger}_{r_{1}}a_{s_{1}}|\Phi(q)\rangle}{\langle\Phi(p)|\Phi(q)\rangle}\;,$ (44a)
$\kappa_{s_{1}s_{2}}(p,q)$ $\equiv\frac{\langle\Phi(p)|a_{s_{2}}a_{s_{1}}|\Phi(q)\rangle}{\langle\Phi(p)|\Phi(q)\rangle}\;,$ (44b)
$\bar{\kappa}^{*}_{r_{1}r_{2}}(p,q)$ $\equiv\frac{\langle\Phi(p)|a^{\dagger}_{r_{1}}a^{\dagger}_{r_{2}}|\Phi(q)\rangle}{\langle\Phi(p)|\Phi(q)\rangle}\;.$ (44c)
@@
$\begin{split}\langle\Phi(p)|\mathcal{E}^{(2)\,2\text{b}}_{\ell}|\Phi(q)\rangle&=\frac{1}{4}\sum_{ijkl}(\bar{\mathcal{\epsilon}}^{(2)}_{\ell})_{ijkl}[\rho_{ki}(p,q)\rho_{lj}(p,q)-\rho_{kj}(p,q)\rho_{li}(p,q)+\bar{\kappa}^{*}_{ij}(p,q)\kappa_{kl}(p,q)]\\ &=\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{\ell}\rho(p,q)]\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{-\ell}\rho(p,q)]-\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{\ell}\rho(p,q)\mathcal{\epsilon}^{(1)}_{-\ell}\rho(p,q)]\\ &\quad+\frac{1}{2}\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{\ell}\kappa(p,q)(\mathcal{\epsilon}^{(1)}_{-\ell})^{T}\bar{\kappa}^{\dagger}(p,q)]-\frac{1}{2}\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{\ell}\kappa^{T}(p,q)(\mathcal{\epsilon}^{(1)}_{-\ell})^{T}\bar{\kappa}^{\dagger}(p,q)]\\ &=\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{\ell}\rho(p,q)]\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{-\ell}\rho(p,q)]-\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{\ell}\rho(p,q)\mathcal{\epsilon}^{(1)}_{-\ell}\rho(p,q)]+\mathrm{Tr}[\mathcal{\epsilon}^{(1)}_{\ell}\kappa(p,q)(\mathcal{\epsilon}^{(1)}_{-\ell})^{T}\bar{\kappa}^{\dagger}(p,q)]\end{split}$ (46)
where advantage was taken of the separable character of the two-body matrix elements (Eq. (29b)) as well as of the antisymmetry of the density $\kappa$ to obtain an expression that is much less costly to compute than for a general two-body operator.
II.8 Classical rigid rotor interpretation
The interpretation of the normalized mean-squared eccentricity $\langle\delta\epsilon^{2}_{\ell}\rangle_{0^{+}}$ within the classical rigid-rotor (RR) model allows for insightful conclusions regarding the nature of multi-particle correlations [15].
@@ II.8 Classical rigid rotor interpretation
$\rho^{(1)}_{\mathrm{intr}}(\mathbf{r})$ $\equiv A\frac{e^{-r^{2}/2R(\theta)^{2}}}{(2\pi)^{3/2}R^{3}}\;,$ (47)
@@
$R(\theta)\equiv R[1+\beta_{20}Y_{2}^{0}(\theta)+\beta_{30}Y_{3}^{0}(\theta)],$ (48)
and where $\theta$ is the angle in the intrinsic frame with respect to the symmetry axis.
@@
$\rho^{(1)}_{\mathrm{intr}}(\mathbf{r})$ $\approx A\frac{e^{-r^{2}/2R^{2}}}{(2\pi)^{3/2}R^{3}}[1+\frac{r^{2}}{R^{2}}\beta_{20}Y_{2}^{0}(\theta)+\frac{r^{2}}{R^{2}}\beta_{30}Y_{3}^{0}(\theta)]\;,$ (49)
@@
$\rho^{(1)}_{\mathrm{lab}}(\mathbf{r})\equiv\frac{1}{4\pi}\int_{\Omega_{\mathrm{or}}}\rho^{(1)}_{\mathrm{intr}}(R_{\Omega_{\rm or}}^{-1}\mathbf{r})=A\frac{e^{-r^{2}/2R^{2}}}{(2\pi)^{3/2}R^{3}}\;,$ (50)
@@
Calculating the expectation value of the two-body part of the squared eccentricity operator requires the local two-body density, which is obtained in the classical RR model according to [^4] Because of its classical character, the RR model only consider the direct part of the two-body density, which is thus not antisymmetric under the exchange of the two particles coordinates. As a result, its integration over space is equal to $A^{2}$ instead of A(A1)A(A-1) as it would be the case for a quantal two-body density.  [^5] The fact that $\rho^{(2)}_{\mathrm{lab}}(\mathbf{r}_{1},\mathbf{r}_{2})$ has a contribution of the type $\beta_{\ell 0}^{2}r_{1}^{\ell}r_{2}^{\ell}\cos(\ell(\phi_{1}-\phi_{2}))$ was argued for the first time in Ref. [9] for a number of test cases.
$\rho^{(2)}_{\mathrm{lab}}(\mathbf{r}_{1},\mathbf{r}_{2})$ $\equiv\frac{1}{4\pi}\int_{\Omega_{\mathrm{or}}}\rho^{(1)}_{\mathrm{intr}}(R_{\Omega_{\rm or}}^{-1}\mathbf{r}_{1})\rho^{(1)}_{\mathrm{intr}}(R_{\Omega_{\rm or}}^{-1}\mathbf{r}_{2})$
$=A^{2}\frac{e^{-(r^{2}_{1}+r^{2}_{2})/2R^{2}}}{32\pi^{4}R^{6}}\times[4\pi$ (51)
@@
Evaluating Eq. (21) using Eqs. (50) and (II.8) leads to
$\langle\delta\epsilon^{2}_{2}\rangle^{\text{RR}}_{0^{+}_{1}}=\frac{1}{A}+\frac{3}{4\pi}\beta^{2}_{20}\;,\quad\langle\delta\epsilon^{2}_{3}\rangle^{\text{RR}}_{0^{+}_{1}}=\frac{16}{3\pi A}+\frac{2048}{245\pi^{3}}\beta^{2}_{30}\;.$ (53)
@@
$\mathcal{B}^{2}_{\ell}(\text{HE})\equiv\frac{4\pi}{3}\langle\delta\epsilon^{2}_{\ell}\rangle^{2\text{b}}\;.$ (54)
III Kumar operator
III.1 Definition